OpenStudy (anonymous):

The circumference of a circle with a diameter of 3.73 in is _____in. Round answer to nearest hundredth.

OpenStudy (aravindg):

\[Circumference =2 \times \pi \times r \]

OpenStudy (aravindg):

where r is radius of the circle

OpenStudy (aravindg):

can you do it now @flyboy13th